Two different fruit juices are mixed together. 7 liters of the first fruit juice contains 60% apple juice, and 5 liters of the second fruit juice contains 30% apple juice. Find the percentage of the new mix that is apple juice. 



1
Expert's answer
2021-08-16T09:50:53-0400

7 liters of the first fruit juice contain "7\cdot\dfrac{60\%}{100\%}=4.2" liters of the apple juice.


5 liters of the second fruit juice contain "5\cdot\dfrac{30\%}{100\%}=1.5" liters of the apple juice.


"5+7=12" liters of the new mix contain "4.2+1.5=5.7" liters of the apple juice.


Then

"\dfrac{5.7}{12}\cdot100\%=47.5\%"

New mix contains 47.5 % apple juice.